# DiffAugment: Diffusion based Long-Tailed Visual Relationship Recognition

### 01 Jan 2024 v1

@@ Abstract
The task of Visual Relationship Recognition (VRR) aims to identify relationships between two interacting objects in an image and is particularly challenging due to the widely-spread and highly imbalanced distribution of $<$subject, relation, object$>$ triplets. To overcome the resultant performance bias in existing VRR approaches, we introduce DiffAugment – a method which first augments the tail classes in the linguistic space by making use of WordNet and then utilizes the generative prowess of Diffusion Models to expand the visual space for minority classes. We propose a novel hardness-aware component in diffusion which is based upon the hardness of each $<$S,R,O$>$ triplet and demonstrate the effectiveness of hardness-aware diffusion in generating visual embeddings for the tail classes. We also propose a novel subject and object based seeding strategy for diffusion sampling which improves the discriminative capability of the generated visual embeddings. Extensive experimentation on the GQA-LT dataset shows favorable gains in the subject/object and relation average per-class accuracy using Diffusion augmented samples.
@@ 4.1 Experimental settings
In this work, we perform experiments on the GQA-LT dataset built on top of Visual Genome by [2]. It has $72,580$ training images, $2,573$ validation images and $7,722$ test images; with $1,703$ objects and $310$ relations. The most frequent object and relation has $374,282$ and $1,692,068$ examples, and the least frequent has $1$ and $2$ respectively. Following the strategy used by [2], the dataset is split into three parts– Many, Medium, Few and the selection ratio of each split is based on the frequency of each class: Many (top $5\%$s - 86 classes for S/O and 16 classes for R), Medium(middle $15\%$ 255 classes for S/O and 46 classes for R), Few (remaining $80\%$ 1362 classes for S/O and 248 classes for R).
@@ 4.4 Qualitative Results
5 Limitations and Future Work
Although Diffusion based augmentation improves the overall performance of any VRR model in long-tail setting, it comes at the cost of reduced performance on head classes - both for subjects/objects as well as relations. This is the case with many imbalance alleviating loss functions as well. Moreover, due to the high training/sampling time and large compute requirements, the number of augmented visual embeddings that have been utilized is much less than the original dataset size ($\approx$100K versus $\approx$3.8M). Therefore, for further exploration, more advanced Diffusion formulations such as Class-balanced Diffusion Models [22] with fast training/sampling strategies can be utilized to observe the effect of Diffusion based Augmentation. Another important direction would be to utilize text-to-image foundation Diffusion Models that have been trained on huge amounts of data [25] to improve the visual relationship understanding of tail classes.

### 01 Mar 2024 v2

@@ Abstract
The task of Visual Relationship Recognition (VRR) aims to identify relationships between two interacting objects in an image and is particularly challenging due to the widely-spread and highly imbalanced distribution of $<$subject, relation, object$>$ triplets. To overcome the resultant performance bias in existing VRR approaches, we introduce DiffAugment – a method which first augments the tail classes in the linguistic space by making use of WordNet and then utilizes the generative prowess of Diffusion Models to expand the visual space for minority classes. We propose a novel hardness-aware component in diffusion which is based upon the hardness of each $<$S,R,O$>$ triplet and demonstrate the effectiveness of hardness-aware diffusion in generating visual embeddings for the tail classes. We also propose a novel subject and object based seeding strategy for diffusion sampling which improves the discriminative capability of the generated visual embeddings. Extensive experimentation on the GQA-LT dataset shows favorable gains in the subject/object and relation average per-class accuracy using Diffusion augmented samples.Our code will be released.
@@ 4.1 Experimental settings
In this work, we perform experiments on the GQA-LT dataset built on top of Visual Genome by [2]. It has $72,580$ training images, $2,573$ validation images and $7,722$ test images; with $1,703$ objects and $310$ relations. The most frequent object and relation has $374,282$ and $1,692,068$ examples, and the least frequent has $1$ and $2$ respectively. Following the strategy used by [2], the dataset is split into three parts– Many, Medium, Few and the selection ratio of each split is based on the frequency of each class: Many (top $5\%$ 86 classes for S/O and 16 classes for R), Medium(middle $15\%$ 255 classes for S/O and 46 classes for R), Few (remaining $80\%$ 1362 classes for S/O and 248 classes for R).